OpenStudy (cutiecomittee123):

The projection of v=(6,4) onto w=(1,1) is _____

OpenStudy (loser66):

you have formula to do it. what is it?

OpenStudy (cutiecomittee123):

p=(v(w))/(absolute value of w)

OpenStudy (cutiecomittee123):

I get (10)/(absolute value of (1,1))

OpenStudy (loser66):

\(proj_a u =\dfrac{u\bullet a}{||a||^2}a\)

OpenStudy (loser66):

your a is w, your u is v. Apply it. show your work, please

OpenStudy (xapproachesinfinity):

Proj of v onto w=v.w/||w|| times w

OpenStudy (loser66):

hey kid, the denominator is power of 2.

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

yes old man just forget it :)

OpenStudy (loser66):

I am busy now, but you must teach me how to do that grid later, kid?

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

just a visualization of what is going on

OpenStudy (xapproachesinfinity):

at any rate do you know how to use the formula

OpenStudy (cutiecomittee123):

sorta I am confused on how to calculate the absolute value of a point though

OpenStudy (cutiecomittee123):

@xapproachesinfinity

OpenStudy (cutiecomittee123):

So then wouldnt it be 10? because after plugging in the numbers I get 10/1 which is just 10

OpenStudy (cutiecomittee123):

@Loser66

OpenStudy (phi):

|a| can be thought of as the length of a vector from (0,0) to the point a use pythagoras. e.g. if a=(x,y) then |a| = \(\sqrt{x^2+y^2}\)

OpenStudy (cutiecomittee123):

Okay

OpenStudy (cutiecomittee123):

explain more because that kinda confused me